OpenStudy (abdumu16):

Your pension plan is an annuity with a guaranteed return of 4% per year (compounded quarterly). You can afford to put $1,100 per quarter into the fund, and you will work for 40 years before retiring. After you retire, you will be paid a quarterly pension based on a 25-year payout. How much will you receive each quarter? (Round your answer to the nearest cent.)

OpenStudy (will.h):

Cool question

OpenStudy (will.h):

1st you'll have to find the total retirement returnings after the interest rate using P(1 + r)^x R is rate of change X is time in years P is the principal P is 1100 per quarter which means per entire year it would be 4400 X is 40 years R is 4% which means 0.04 4400(1+0.04)^40 Total retirement returnings= 21124.491 Now you'll be paid on 25 years quarterly Divide that by 25 years to see tha amount annually. 21124.491/25 = 844.97964 Again divide that by 4 because the pension will be given quarterly 844.97964/4 = 211.24491 dollars